OpenStudy (anonymous):

Write the equation of the tangent line of y=2cosx at x=-pi/2??

hartnn (hartnn):

to get equation of line, you'll need slope of tangent slope of tangent at x = -pi/2 = y' at x = (-pi/2) so first can you find the derivative of y ?

OpenStudy (anonymous):

is the dervative of y=-sinx?

hartnn (hartnn):

you mean \( y' = -2\sin x ?\)

hartnn (hartnn):

then you just plug in x = -p/2 to get the slope \(m= -2 \sin (\pi/2 )\)